OpenStudy (anonymous):

one year after the celebrity first began posting comments on a website, she had 23,100 followers.Five years after this celebrity first began posting comments,she had 165,100 followers. (A) find the slope of the line that passes through the points (1,2300), and (5,165100). (B) interpret the slope as a rate of change. Thus the slope is?

OpenStudy (kropot72):

The slope m is found from the formula \[m=\frac{y _{1}-y _{2}}{x _{1}-x _{2}}\]

OpenStudy (anonymous):

What would be the answer?

OpenStudy (kropot72):

\[m=\frac{2300-165100}{1-5}=you\ can\ calculate\]

OpenStudy (anonymous):

does that mean i would take 2300-165100=162,800?

OpenStudy (kropot72):

Nearly right. 2300 - 165100 = -162800

OpenStudy (kropot72):

And 1 - 5 = -4 \[\frac{-162800}{-4}=?\]

OpenStudy (anonymous):

162,796?

OpenStudy (kropot72):

The fraction can be made to look simpler by multiplying both the numerator and the denominator by -1: \[m=\frac{162800}{4}=?\]

OpenStudy (anonymous):

40,700?

OpenStudy (kropot72):

Correct! The slope of 40,700 means that the celebrity is gaining followers at the rate of 40,700 per year.

OpenStudy (anonymous):

it states that its not correct

OpenStudy (kropot72):

When I check your question there is an incorrect ordered pair. The first ordered pair should read (1, 23100). Therefore the slope needs to be recalculated as \[m=\frac{23100-165100}{1-5}=\frac{-142000}{-4}=35500\]

OpenStudy (kropot72):

That means the celebrity is gaining followers at the rate of 35500 per year.

OpenStudy (anonymous):

thats correct 35500